6.8 g H2O2 is dissolved in 0.2 L solution. Then solution will be labelled as

Grade:12th pass

2 Answers

Sandeep Bhamoo
5 years ago
Since mol. Mass of H2O2=34 Hence mole of H2O2=6.8÷34=1÷5=0.2. AND, H2O2-------> H2O + 1÷2(O2) since one mole H2O2 gives 1÷2 mole O2 hence 0.2 mole H2O2 will give 0.1 mole O2 ,that is equal to 2.24 l of O2 in a sol. of 0.2 l solution,that is 0.2l sol. gives2.24l O2 hence 1l sol. would give 11.2l O2 . Hence it would be labelled as "11.2 V H2O2"
Sujit Kumar
5 years ago
The answer is as follows...
 
Given\ mass = 6.8g
Gram\ molecular\ mass\ (of\ H_{2}O_{2})=2(1)=2(16)=34
No.\ of\ moles = \frac{Given\ mass}{Gram\ molecular\ mass}
\rightarrow No.\ of\ moles\ = \frac{6.8}{34}=0.2
Molarity=\frac{No.\ of\ moles\ of\ solute}{Volume\ of solution\ (in L)}
\rightarrow Molarity=\frac{0.2}{0.2}=1M
 
Answer \mapsto 1M
 
(Didn’t realize the answer needs some text... so im just filling up the space for this to get submitted! Hope you understand the solution given above!)